# On multi-graded Proj schemes

### v4

@@ 3.5 Relative Proj schemes
Construction 3.22 ((Relative Proj)).
@@ 3.6 Proj schemes, blowups and dilatations
Let $A$ be a ring and $\{J_{i}:i\in I\}$ be a family of ideals of $A$ parametrized by a finite set $I$. Recall the ring ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A=\bigoplus_{\nu\in\mathbb{N}^{I}}J_{i}^{\nu}$ (see Example 2.13 and [Ma23b, Definition 2.28]), which is $\mathbb{Z}^{I}$-graded. Any family $\{a_{i}:i\in I\}$ of elements of $A$ such that $a_{i}\in J_{i}$ for each $i\in I$ can be considered as a relevant family in the $\mathbb{Z}^{I}$-graded ring ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A$, so that we have an associated affine open subscheme
@@
The goal of the present subsection is to study related connections and compatibilities between dilatations and these specific Proj schemes, and to show that this allows to easily provide a proof of the existence of (projective) blowups, possibly with multiple centers. This section expands [Ma23b, § 3.8]. We start with a multi-centered version of [StP, Tag 0804].
@@
be the associated Rees algebra (cf. e.g. [Ma23b, §3.1]). This is canonically a $\mathbb{Z}^{I}$-graded quasi-coherent $\mathcal{O}_{X}$-algebra. Using the construction of §3.5 we now define an $X$-scheme
$\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X:={\operatorname{Proj}}^{\mathbb{Z}^{I}}_{X}(\mathrm{Bl}_{\{\mathcal{J}_{i}:i\in I\}}\mathcal{O}_{X}).$
@@
The canonical morphism $\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X\to X$ is separated.
@@
Since the property of being separated is local on the target (see e.g. [StP, Tag 02KU]), we can assume $X={\operatorname{Spec}}(A)$ is affine, and by [StP, Tag 01KV] the claim will follow if in this case we prove that $\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X$ is a separated scheme. We proceed with the usual local notations from the beginning of the subsection. Consider the collection of families $\{a_{i}:i\in I\}$ with $a_{i}\in J_{i}$ for all $i\in I$, and let $\mathcal{F}\subset\mathcal{F}_{A}$ be the subset consisting of the multiplicative subsets generated by such families. By Lemma 3.23, we have ${\operatorname{Proj}}^{\mathbb{Z}^{I}}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)={\operatorname{Proj}}^{\mathbb{Z}^{I}}_{{\mathcal{F}}}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)$. Using the notation from Proposition 3.7, for any $T\in{\mathcal{F}}$ we have $C_{T}=(\mathbb{R}_{\geqslant 0})^{I}\subset\mathbb{R}^{I}$. So if $T,T^{\prime}$ belong to ${\mathcal{F}}$, $C_{T}\cap C_{T^{\prime}}$ has nonempty interior. Now Proposition 3.7 finishes the proof. ∎
@@
If $C$ is an effective Cartier divisor on $X$, then $\left(\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X\right)\times_{X}C$ is an effective Cartier divisor on $\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X$.
@@
This statement is local on $X$, so we can assume $X={\operatorname{Spec}}(A)$ for some ring $A$, and write $J_{i}$ for the global sections of ${\mathcal{J}}_{i}$. In view of Lemma 3.23, in this case it suffices to prove that for any family $\{a_{i}:i\in I\}$ with $a_{i}\in J_{i}$ for any $i$, if $C$ is an effective Cartier divisor on $X$ then ${\operatorname{Spec}}(A[\{\frac{J_{i}}{a_{i}}:i\in I\}])\times_{X}C$ is an effective Cartier divisor on ${\operatorname{Spec}}(A[\{\frac{J_{i}}{a_{i}}:i\in I\}])$. This statement is a special case of [Ma23b, Proposition 3.18]. ∎
In the following statement we use the notation of dilatations of schemes from [Ma23b] (see also [MRR23, DMdS23]). From now on, for simplicity of notation we set
$X^{\prime}:=\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X,\quad Y_{i}^{\prime}:=Y_{i}\times_{X}X^{\prime}.$
@@
For all $i\in I$, $Y_{i}^{\prime}$ is an effective Cartier divisor on $X^{\prime}$. Moreover, the composition $\mathrm{Bl}^{\{Y_{i}^{\prime}:i\in I\}}_{\{\varnothing:i\in I\}}X^{\prime}\to X^{\prime}\to X$ is an open immersion; in fact, this is just the inclusion $X\smallsetminus(\cup_{i\in I}Y_{i})\to X$.
@@
This statement is local on $X$, so we can assume $X={\operatorname{Spec}}(A)$ for some ring $A$, and write $J_{i}$ for the global sections of ${\mathcal{J}}_{i}$. In this case, by Lemma 3.23 the scheme $X^{\prime}=\underline{\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X$ is covered by affine open subschemes ${\operatorname{Spec}}\big{(}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)_{(\{a_{i}:i\in I\})}\big{)}$ where $\{a_{i}:i\in I\}$ is a family of elements of $A$ such that $a_{i}\in J_{i}$ for any $i$. As explained above, $({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)_{(\{a_{i}:i\in I\})}$ is the dilatation ring $A[\{\frac{J_{i}}{a_{i}}:i\in I\}]$. Now the first assertion follows from [Ma23b, Facts 2.10 & 2.25]. The second assertion follows from the formula
$\big{(}A[\big{\{}\frac{J_{i}}{a_{i}}:i\in I\big{\}}]\big{)}\Big{[}\Big{\{}\frac{A[\{\frac{J_{i}}{a_{i}}:i\in I\}]}{a_{i}}:i\in I\Big{\}}\Big{]}=A[\big{\{}\frac{A}{a_{i}}:i\in I\big{\}}]$
(see e.g. the computation in [Ma23b, Proof of Corollary 2.23]) and the fact the the right-hand side is the localization of $A$ with respect to the $a_{i}$’s, see [Ma23b, Fact 2.11]. ∎
@@
First we show that for $T$ and $\pi$ as in the statement, there exists at most one morphism of $X$-schemes $T\to X^{\prime}$. In fact, let $\phi,\phi^{\prime}:T\to X^{\prime}$ be two morphisms of $X$-schemes. Denote by $U$ the open complement of the sum of the effective Cartier divisors $T\times_{X}Y_{i}$; then $U$ is scheme theoretically dense in $T$ by [StP, Tag 07ZU]. By Lemma 3.25, the morphism $X^{\prime}\to X$ is separated; as a consequence, the difference kernel $\mathrm{ker}(\phi,\phi^{\prime})$ (see [GW20, Definition 9.1]) is a closed subscheme of $T$ by [GW20, Definition/Proposition 9.7]. Since $U\times_{X}Y_{i}=\varnothing$ for all $i$, the universal property of dilatations (see [Ma23b, Proposition 3.17]) implies that both $\phi_{|U}$ and $\phi^{\prime}_{|U}$ factor through $\mathrm{Bl}_{\{\varnothing:i\in I\}}^{\{Y_{i}^{\prime}:i\in I\}}X^{\prime}$. By Proposition 3.27, $\mathrm{Bl}_{\{\varnothing:i\in I\}}^{\{Y_{i}^{\prime}:i\in I\}}X^{\prime}$ is an open subscheme of $X$, so that this implies that $\phi_{|U}=\phi^{\prime}_{|U}$. We deduce that $U\subset\mathrm{ker}(\phi,\phi^{\prime})$, and then that $\mathrm{ker}(\phi,\phi^{\prime})=T$, so that $\phi=\phi^{\prime}$.
To prove existence we first consider the following setting. We assume that $X={\operatorname{Spec}}(A)$ and $T={\operatorname{Spec}}(R)$ for some rings $A$ and $R$, and denote by $J_{i}$ the global sections of ${\mathcal{J}}_{i}$, so that $Y_{i}={\operatorname{Spec}}(A/J_{i})$, and by $f:A\to R$ the morphism of rings corresponding to $\pi:T\to X$. With this notation, we further assume that, for any $i\in I$, we have $f(J_{i})R=\gamma_{i}R$ for some non-zero-divisor $\gamma_{i}\in R$.
@@
where $(f(J)R)^{\nu}$ denotes the ideal $\prod_{i\in I}(f(J_{i})R)^{\nu_{i}}\subset R$. Now since $f(J_{i})R=\gamma_{i}R$ with $\gamma_{i}$ non-zero-divisor for any $i$, we have a canonical isomorphism of $R$-algebras $\bigoplus_{\nu\in\mathbb{N}^{I}}(f(J)R)^{\nu}\cong R[\mathbb{N}^{I}]$. So we get a canonical morphism of graded $A$-algebras
@@
Now we prove the existence of a morphism of $X$-schemes $T\to X^{\prime}$ in general. For this we choose an affine open covering $X=\bigcup_{\alpha\in A}X_{\alpha}$. Then for any $\alpha\in A$ and $i\in I$, $(T\times_{X}X_{\alpha})\times_{X_{\alpha}}(Y_{i}\times_{X}X_{\alpha})$ is an effective Cartier divisor on $T\times_{X}X_{\alpha}$. Hence there exists an affine open covering $T\times_{X}X_{\alpha}=\bigcup_{\beta\in B_{\alpha}}T_{\alpha,\beta}$ such that each $T_{\alpha,\beta}\times_{X_{\alpha}}(Y_{i}\times_{X}X_{\alpha})$ is the closed subscheme defined by a non-zero-divisor in $\mathcal{O}(T_{\alpha,\beta})$, see e.g. [StP, Tag 01WS]. By the case treated above, for any $\alpha,\beta$ there exists a morphism of $X_{\alpha}$-schemes $T_{\alpha,\beta}\to X^{\prime}\times_{X}X_{\alpha}$. By uniqueness, these morphisms coincide on the intersections of the $T_{\alpha,\beta}$’s, hence they glue to a morphism of $X$-schemes $T\to X^{\prime}$, as desired.
@@
In other words, Proposition 3.28 says that $\underline{{\operatorname{Bl}}}_{\{Y_{i}:i\in I\}}X$ is a final object in the category ${\operatorname{Sch}}^{\{Y_{i}:i\in I\}\text{-reg}}_{X}$ of [Ma23b, Definition 3.4].
We finish this subsection with some applications of Proposition 3.28. For the first one, recall the “$+$” operation on closed subschemes of $X$, see [Ma23b, Notation 3.1].
@@
We have a canonical isomorphism of $X$-schemes $\underline{{\operatorname{Bl}}}_{\{Y_{i}:i\in I\}}X\cong\underline{{\operatorname{Bl}}}_{\sum_{i\in I}Y_{i}}X$.
@@
It is enough to prove that $\underline{{\operatorname{Bl}}}_{\sum_{i\in I}Y_{i}}X$ is also a blowup of $X$ with center $\{Y_{i}:i\in I\}$. This is immediate using that $\underline{{\operatorname{Bl}}}_{\sum_{i\in I}Y_{i}}X$ is a blowup of $X$ with center $\{\sum_{i\in I}Y_{i}\}$ and the fact that effective Cartier divisors form a face of the monoid of closed subschemes (see [Ma23b, Notation 3.1]). ∎
@@
$\underline{{\operatorname{Bl}}}_{\{Y_{i}:i\in I\}}X\cong\underline{{\operatorname{Bl}}}_{\{Y_{j}\times_{X}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X:j\in J\}}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X.$
@@
It is enough to prove that $\underline{{\operatorname{Bl}}}_{\{Y_{j}\times_{X}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X:j\in J\}}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X$ is also a blowup of $X$ with center $\{Y_{i}:i\in I\}$. Using Proposition 3.26, we see that $\underline{{\operatorname{Bl}}}_{\{Y_{j}\times_{X}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X:j\in J\}}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X$ belongs to ${\operatorname{Sch}}_{X}^{\{Y_{i}:i\in I\}\text{-reg}}$. Let $T\to X$ be in ${\operatorname{Sch}}_{X}^{\{Y_{i}:i\in I\}\text{-reg}}$. Applying twice Proposition 3.28, we get a unique $X$-morphism $T\to\underline{{\operatorname{Bl}}}_{\{Y_{j}\times_{X}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X:j\in J\}}\underline{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X$. This finishes the proof. ∎

### v5

@@ 3.5 Relative Proj schemes
Construction 3.24 ((Relative Proj)).
@@ 3.6 Proj schemes, blowups and dilatations
Let $A$ be a ring and $\{J_{i}:i\in I\}$ be a family of ideals of $A$ parametrized by a finite set $I$. Recall the ring ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A=\bigoplus_{\nu\in\mathbb{N}^{I}}J_{i}^{\nu}$ (see Example 2.14 and [Ma24, Definition 2.34]), which is $\mathbb{Z}^{I}$-graded. Any family $\{a_{i}:i\in I\}$ of elements of $A$ such that $a_{i}\in J_{i}$ for each $i\in I$ can be considered as a relevant family in the $\mathbb{Z}^{I}$-graded ring ${\operatorname{Bl}}_{\{J_{i}:i\in I\}}A$, so that we have an associated affine open subscheme
@@
The goal of the present subsection is to study related connections and compatibilities between dilatations and these specific Proj schemes, and to show that this allows to easily provide a proof of the existence of (projective) blowups, possibly with multiple centers. This section was announced in [Ma24, § 3.9]. We start with a multi-centered version of [StP, Tag 0804].
@@
be the associated Rees algebra (cf. e.g. [Ma24, §3.1]). This is canonically a $\mathbb{Z}^{I}$-graded quasi-coherent $\mathcal{O}_{X}$-algebra. Using the construction of §3.5 we now define an $X$-scheme
${\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X:={\operatorname{Proj}}^{\mathbb{Z}^{I}}_{X}(\mathrm{Bl}_{\{\mathcal{J}_{i}:i\in I\}}\mathcal{O}_{X}).$
@@
The canonical morphism ${\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X\longrightarrow X$ is separated.
@@
Since the property of being separated is local on the target (see e.g. [StP, Tag 02KU]), we can assume $X={\operatorname{Spec}}(A)$ is affine, and by [StP, Tag 01KV] the claim will follow if in this case we prove that ${\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X$ is a separated scheme. We proceed with the usual local notations from the beginning of the subsection. Consider the collection of families $\{a_{i}:i\in I\}$ with $a_{i}\in J_{i}$ for all $i\in I$, and let $\mathcal{F}\subset\mathcal{F}_{A}$ be the subset consisting of the multiplicative subsets generated by such families. By Lemma 3.25, we have ${\operatorname{Proj}}^{\mathbb{Z}^{I}}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)={\operatorname{Proj}}^{\mathbb{Z}^{I}}_{{\mathcal{F}}}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)$. Using the notation from Proposition 3.8, for any $T\in{\mathcal{F}}$ we have $C_{T}=(\mathbb{R}_{\geqslant 0})^{I}\subset\mathbb{R}^{I}$. So if $T,T^{\prime}$ belong to ${\mathcal{F}}$, $C_{T}\cap C_{T^{\prime}}$ has nonempty interior. Now Proposition 3.8 finishes the proof. ∎
@@
If $C$ is an effective Cartier divisor on $X$, then $\left({\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X\right)\times_{X}C$ is an effective Cartier divisor on ${\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X$.
@@
This statement is local on $X$, so we can assume $X={\operatorname{Spec}}(A)$ for some ring $A$, and write $J_{i}$ for the global sections of ${\mathcal{J}}_{i}$.
In view of Lemma 3.25, in this case it suffices to prove that for any family $\{a_{i}:i\in I\}$ with $a_{i}\in J_{i}$ for any $i$, if $C$ is an effective Cartier divisor on $X$ then ${\operatorname{Spec}}(A[\{\frac{J_{i}}{a_{i}}:i\in I\}])\times_{X}C$ is an effective Cartier divisor on ${\operatorname{Spec}}(A[\{\frac{J_{i}}{a_{i}}:i\in I\}])$. This statement is a special case of [Ma24, Proposition 3.18]. ∎
In the following statement we use the notation of dilatations of schemes from [Ma24] (see also [MRR23, DMdS23]). From now on, for simplicity of notation we set
$X^{\prime}:={\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X,\quad Y_{i}^{\prime}:=Y_{i}\times_{X}X^{\prime}.$
@@
For all $i\in I$, $Y_{i}^{\prime}$ is an effective Cartier divisor on $X^{\prime}$. Moreover, the composition $\mathrm{Bl}^{\{Y_{i}^{\prime}:i\in I\}}_{\{\varnothing:i\in I\}}X^{\prime}\to X^{\prime}\to X$ is an open immersion [^4] Recall that if $X$ is a scheme, $\varnothing$ denotes the empty closed subscheme corresponding to the quasi-coherent ideal ${\mathcal{O}}_{X}$. In particular $\mathrm{Bl}^{\{Y_{i}^{\prime}:i\in I\}}_{\{\varnothing:i\in I\}}X^{\prime}$ denotes the dilatation of $X^{\prime}$ relatively to the multi-center $\{[\varnothing,Y_{i}^{\prime}]:i\in I\}$ ($Y_{i}^{\prime}$ is a Cartier divisor in $X^{\prime}$ by assumption). ; in fact, this is just the inclusion $X\smallsetminus(\cup_{i\in I}Y_{i})\to X$.
@@
This statement is local on $X$, so we can assume $X={\operatorname{Spec}}(A)$ for some ring $A$, and write $J_{i}$ for the global sections of ${\mathcal{J}}_{i}$. In this case, by Lemma 3.25 the scheme $X^{\prime}={\mathrm{Bl}}_{\{Y_{i}:i\in I\}}X$ is covered by affine open subschemes ${\operatorname{Spec}}\big{(}({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)_{(\{a_{i}:i\in I\})}\big{)}$ where $\{a_{i}:i\in I\}$ is a family of elements of $A$ such that $a_{i}\in J_{i}$ for any $i$. As explained above, $({\operatorname{Bl}}_{\{J_{i}:i\in I\}}A)_{(\{a_{i}:i\in I\})}$ is the dilatation ring $A[\{\frac{J_{i}}{a_{i}}:i\in I\}]$. Now the first assertion follows from [Ma24, Facts 2.10 & 2.28]. The second assertion follows from the equalities
$\big{(}A[\big{\{}\frac{J_{i}}{a_{i}}:i\in I\big{\}}]\big{)}\Big{[}\Big{\{}\frac{A[\{\frac{J_{i}}{a_{i}}:i\in I\}]}{a_{i}}:i\in I\Big{\}}\Big{]}$ $=\big{(}A[\big{\{}\frac{J_{i}}{a_{i}}:i\in I\big{\}}]\big{)}\Big{[}\Big{\{}\frac{A}{a_{i}}:i\in I\Big{\}}\Big{]}$
$=A[\big{\{}\frac{J_{i}}{a_{i}}:i\in I\big{\}}\sqcup\{\frac{A}{a_{i}}:i\in I\big{\}}]$
$=A[\big{\{}\frac{A}{a_{i}}:i\in I\big{\}}]$
$=A_{\{a_{i}\}_{i\in I}}$
(use e.g. [Ma24, 2.22, 2.28, 2.11]) and the fact that $A_{\{a_{i}\}_{i\in I}}$ is the localization of $A$ with respect to the $a_{i}$’s. ∎
@@
First we show that for $T$ and $\pi$ as in the statement, there exists at most one morphism of $X$-schemes $T\to X^{\prime}$. In fact, let $\phi,\phi^{\prime}:T\to X^{\prime}$ be two morphisms of $X$-schemes. Denote by $U$ the open complement of the sum of the effective Cartier divisors $T\times_{X}Y_{i}$; then $U$ is scheme theoretically dense in $T$ by [StP, Tag 07ZU]. By Lemma 3.27, the morphism $X^{\prime}\to X$ is separated; as a consequence, the difference kernel $\mathrm{ker}(\phi,\phi^{\prime})$ (see [GW20, Definition 9.1]) is a closed subscheme of $T$ by [GW20, Definition/Proposition 9.7]. Since $U\times_{X}Y_{i}=\varnothing$ for all $i$, the universal property of dilatations (see [Ma24, Proposition 3.17]) implies that both $\phi_{|U}$ and $\phi^{\prime}_{|U}$ factor through $\mathrm{Bl}_{\{\varnothing:i\in I\}}^{\{Y_{i}^{\prime}:i\in I\}}X^{\prime}$. By Proposition 3.29, $\mathrm{Bl}_{\{\varnothing:i\in I\}}^{\{Y_{i}^{\prime}:i\in I\}}X^{\prime}$ is an open subscheme of $X$, so that this implies that $\phi_{|U}=\phi^{\prime}_{|U}$. We deduce that $U\subset\mathrm{ker}(\phi,\phi^{\prime})$, and then that $\mathrm{ker}(\phi,\phi^{\prime})=T$, so that $\phi=\phi^{\prime}$.
To prove existence we first consider the following setting. We assume that $X={\operatorname{Spec}}(A)$ and $T={\operatorname{Spec}}(R)$ for some rings $A$ and $R$, and denote by $J_{i}$ the global sections of ${\mathcal{J}}_{i}$, so that $Y_{i}={\operatorname{Spec}}(A/J_{i})$, and by $f:A\to R$ the homomorphism of rings corresponding to $\pi:T\to X$. With this notation, we further assume that, for any $i\in I$, we have $f(J_{i})R=\gamma_{i}R$ for some non-zero-divisor $\gamma_{i}\in R$.
@@
where $(f(J)R)^{\nu}$ denotes the ideal $\prod_{i\in I}(f(J_{i})R)^{\nu_{i}}\subset R$. Now since $f(J_{i})R=\gamma_{i}R$ with $\gamma_{i}$ non-zero-divisor for any $i$, we have a canonical isomorphism of $R$-algebras $\bigoplus_{\nu\in\mathbb{N}^{I}}(f(J)R)^{\nu}\cong R[\mathbb{N}^{I}]$. So we get a canonical homomorphism of graded $A$-algebras
@@
Now we prove the existence of a morphism of $X$-schemes $T\to X^{\prime}$ in general. For this we choose an affine open covering $X=\bigcup_{\alpha\in\mathscr{C}}X_{\alpha}$. Then for any $\alpha\in\mathscr{C}$ and $i\in I$, $(T\times_{X}X_{\alpha})\times_{X_{\alpha}}(Y_{i}\times_{X}X_{\alpha})$ is an effective Cartier divisor on $T\times_{X}X_{\alpha}$. Hence there exists an affine open covering $T\times_{X}X_{\alpha}=\bigcup_{\beta\in B_{\alpha}}T_{\alpha,\beta}$ such that each $T_{\alpha,\beta}\times_{X_{\alpha}}(Y_{i}\times_{X}X_{\alpha})$ is the closed subscheme defined by a non-zero-divisor in $\mathcal{O}(T_{\alpha,\beta})$, see e.g. [StP, Tag 01WS]. By the case treated above, for any $\alpha,\beta$ there exists a morphism of $X_{\alpha}$-schemes $T_{\alpha,\beta}\to X^{\prime}\times_{X}X_{\alpha}$. By uniqueness, these morphisms coincide on the intersections of the $T_{\alpha,\beta}$’s, hence they glue to a morphism of $X$-schemes $T\to X^{\prime}$, as desired.
@@
In other words, Proposition 3.30 says that ${{\operatorname{Bl}}}_{\{Y_{i}:i\in I\}}X$ is a final object in the category ${\operatorname{Sch}}^{\{Y_{i}:i\in I\}\text{-reg}}_{X}$ defined in [Ma24, Definition 3.4].
We finish this subsection with some applications of Proposition 3.30. For the first one, recall the “$+$” operation on closed subschemes of $X$, see [Ma24, Notation 3.1].
@@
We have a canonical isomorphism of $X$-schemes ${{\operatorname{Bl}}}_{\{Y_{i}:i\in I\}}X\cong{{\operatorname{Bl}}}_{\sum_{i\in I}Y_{i}}X$.
@@
It is enough to prove that ${{\operatorname{Bl}}}_{\sum_{i\in I}Y_{i}}X$ is also a blowup of $X$ with center $\{Y_{i}:i\in I\}$. This is immediate using that ${{\operatorname{Bl}}}_{\sum_{i\in I}Y_{i}}X$ is a blowup of $X$ with center $\{\sum_{i\in I}Y_{i}\}$ and the fact that effective Cartier divisors form a face of the monoid of closed subschemes (see [Ma24, Notation 3.1]). ∎
@@
${{\operatorname{Bl}}}_{\{Y_{i}:i\in I\}}X\cong{{\operatorname{Bl}}}_{\{Y_{j}\times_{X}{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X:j\in J\}}{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X.$
@@
It is enough to prove that ${{\operatorname{Bl}}}_{\{Y_{j}\times_{X}{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X:j\in J\}}{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X$ is also a blowup of $X$ with center $\{Y_{i}:i\in I\}$. Using Proposition 3.28, we see that ${{\operatorname{Bl}}}_{\{Y_{j}\times_{X}{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X:j\in J\}}{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X$ belongs to ${\operatorname{Sch}}_{X}^{\{Y_{i}:i\in I\}\text{-reg}}$. Let $T\to X$ be in ${\operatorname{Sch}}_{X}^{\{Y_{i}:i\in I\}\text{-reg}}$. Applying twice Proposition 3.30, we get a unique $X$-morphism $T\to{{\operatorname{Bl}}}_{\{Y_{j}\times_{X}{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X:j\in J\}}{{\operatorname{Bl}}}_{\{Y_{k}:k\in K\}}X$. This finishes the proof. ∎